A long solenoid has a diameter of 12.0 cm. When a current i exists in its windings, a uniform magnetic field B = 27.0 mT is produced in its interior. By decreasing i, the field is caused to decrease at the rate of 5.50 mT/s. Calculate the magnitude of the induced electric field at the following distances from the axis of the solenoid.
Answer #1

The induced emf = A*dB/dt = ?*(0.06m)^2*5.50x10^-3T/s = 6.54x10^-5V


Now E*L = emf where L = 2?r


So E at a distance r from the solenoid is 6.54x10^-5/(2?*r)


You didn't specify the distances so just plug in the value of r for each